OpenStudy (anonymous):

the expression 2 log a - (3 log b + log c) is equivalent to

OpenStudy (science0229):

Use these 2 properties of log

OpenStudy (science0229):

\[a \log(x)=\log(a^x)\]

OpenStudy (science0229):

\[\log(a)-\log(b)=\log(\frac{ a }{ b })\]